Find the equations for the horizontal and vertical asymptotes of the following. The answers are the equations of lines, so writing simple values will not suffice.

f(x)=(2x-1)/(x^2+2)
Horizontal:
Vertical:

g(x)=5x/(x-1)
Horizontal:
Vertical:

Answer by stanbon(75887)

Find the equations for the horizontal and vertical asymptotes of the following. The answers are the equations of lines, so writing simple values will not suffice.
f(x)=(2x-1)/(x^2+2)
Horizontal: y=0 because the coefficient of x^2 is 0 in the numerator and 1
in the denominator; so y = 0/1 = 0
----
Vertical: nor vertical asymptotes because x^2+2 can never be 0.
-------------------------------------------------------------------------------
g(x)=5x/(x-1)
Horizontal: y = 5 because the coefficient of x is 5 in the numerator and
1 in the denominator; so y = 5/1 = 5
-----
Vertical: x = 1 because the denominator is zero when x = 1
============================================================================
Cheers,
Stan H.